# XtraLight-MedMamba for Classification of Neoplastic Tubular Adenomas

### v1

@@ 3.1.2 Parallel Vision Mamba Module
The PVM module [26, 29], as shown on the left side of Fig. 4 (part a), mainly consists of Mamba integrated with residual connections, which improve Mamba’s ability to capture deep spatial relationships. Initially, the input $Y_{in}^{C}$ with channel $C$ goes through layer normalization. The features that are fed into Mamba are first branched out as $X_{1}^{C/4}$, $X_{2}^{C/4}$, $X_{3}^{C/4}$ and $X_{4}^{C/4}$, each with channel $C/4$. The Mamba outputs combined with the residual connection from the inputs and adjustment factor for optimization are then concatenated together to obtain the four feature maps. The right side of Fig. 4 (part b) shows the Mamba component used in the PVM layer. These features are concatenated to obtain $Y_{out}^{C}$ with the number of channels as $C$. The concatenated feature outputs are layer normalized and then projected using a Projection operation. By processing deep features in a parallel manner, the PVM module is able to capture multi-scale and intricate features using different kernel sizes. This process significantly reduces the parameters by retaining the same receptive field, which addresses the explosive influence of channel numbers since the number of parameters in Mamba is highly dependent on the number of input channels [29]. The specific operation can be expressed by the following equations:
$X_{C/4}^{1},\;X_{C/4}^{2},\;X_{C/4}^{3},\;X_{C/4}^{4}=S_{p}[\text{LN}(Y_{C}^{out})]$ (1)
$VM\_X_{C/4}^{i}$ $=\text{Mamba}(X_{C/4}^{i})+\theta\cdot X_{C/4}^{i},\quad i=1,2,3,4$ (2)
$Y_{\text{out}}=\text{Cat}\big(VM\_X_{1}^{C/4},\,VM\_X_{2}^{C/4},\,VM\_X_{3}^{C/4},\,VM\_X_{1}^{C/4}\big)$ (3)
$\text{Out}=Pro\big[\text{LN}(Y_{\text{out}})\big]$ (4)
Here, $LN$ is the layer normalization, $S_{p}$ is the split operation, $Mamba$ is the Mamba operation, $\theta$ is the adjustment factor for the residual connection, $Cat$ is the concatenation operation, and $Pro$ is the projection operation. From Eq. 2, we employed parallel Vision Mamba feature processing while keeping the total number of channels constant, allowing us to preserve high accuracy while achieving significant parameter reduction.
@@ 3.1.4 Fixed Non-Negative Orthogonal Classifier for Model Parameter Reduction
The Fixed Non-Negative Orthogonal Classifier or FNOClassifier [8] module is a lightweight neural classifier designed to map feature embeddings into class logits through structured and normalized projections. During initialization, the model partitions the input feature dimensions into approximately equal groups corresponding to the number of output classes. Random permutation ensures that each class is assigned a unique subset of features, with any remaining dimensions distributed randomly among the classes to maintain balance. A binary base matrix is then constructed, where each row indicates the feature subset associated with a class, and is subsequently L2-normalized along the row dimension to enforce orthogonality and numerical stability. The model also includes an optional input feature normalization step where input vectors are scaled by their L2 norm, ensuring consistent magnitude across samples. In the forward pass, the input (normalized or raw) is linearly projected using the scaled base matrix, modulated by a learnable scalar parameter $W$ to produce class scores. This allows the FNOClassifier to serve as an efficient and interpretable projection-based classifier, emphasizing structured feature grouping and normalization for robust classification.
Fixed classifiers that incorporate orthogonality are recognized for their cost efficiency and their capacity to sometimes outperform learnable classifiers in popular benchmarks. However, existing fixed orthogonal classifiers suffer from geometric limitations that fundamentally prevent them from invoking neural collapse [8, 30].
@@
The FNOClassifer was proposed to resolve the issue that previous fixed orthogonal classifiers fail to invoke neural collapse (NC). Neural collapse is a critical phenomenon where last-layer features converge to the simplex ETF (Equiangular Tight Frame) structure necessary for achieving global optimality in a layer-peeled model, due to their inherent geometric limitations [8] [30]. The development of the FNOClassifier relies on an analysis of zero-mean NC that specifically considers orthogonality within non-negative Euclidean space. By satisfying the necessary properties of zero-mean NC, the FNOClassifier is designed to induce the optimal solution and maximize the margin of an orthogonal layer-peeled model. This classifier yields an inherent feature dimension separation effect by mitigating feature interference in continual learning and tackling the limitations of mixup on the hypersphere in imbalanced learning, ultimately demonstrating significant performance improvements across various experiments.
Fixed Non-Negative Orthogonal Classifier (FNOClassifier). Let $C$ be the number of classes and $F$ the feature dimension. We fix a classifier with partial orthogonal weight matrix $W$ as $W\in\mathbb{R}^{F\times C}$ (no learning in $W$) and require
$W^{\top}W$ $=I_{C},\qquad W_{i,j}\geq 0\ \ (\text{element\mbox{-}wise}),$ (5)
where $I_{C}$ is identity matrix and $W_{i,j}$ is the $(i,j)$ element of $W$. The logits which are optionally scaled given by
@@
$\qquad p(c\mid x)=\mathrm{softmax}_{c}\big(z(x)\big).$ (7)
where $z(x)$ is the logit vector, $\gamma$ is a positive scaling parameter controlling the magnitude of the logits and $x$ is the input feature representation. $p(c|x)$ is the probability predicted of an input sample with feature representation where $x$ belongs to class $c$.
Neural collapse. First intra-class collapse occurs. Under zero-mean neural collapse, features from the same class converge toward a shared class mean where $x_{i}^{(c)}$ is the feature output of the $i$-th sample belonging to the class $c$, $n_{c}$ and $\mu_{c}$ are the number of samples and the class mean feature vector for class $c$, respectively.
$\mu_{c}=\frac{1}{n_{c}}\sum_{i=1}^{n_{c}}x_{i}^{(c)}$ (8)
$x_{i}^{(c)}$ $\xrightarrow[]{\text{training}}\mu_{c}\quad\text{(intra-class collapse)}$ (9)
@@
Second, the weighted mean of the class becomes centered and converges to zero [16], [7], [2]. Therefore, the features behave as:
$\sum_{c=1}^{C}\pi_{c}\,\mu_{c}$ $=0\quad\text{(centered class means)}.$ (10)
@@
In (10), $C$ is the total number of classes, the class probability is denoted by $\pi_{c}$ where the global feature distribution is centered at origin and the weighted average of all class means is zero.
The equal-norm constraint is given as:
$|\mu_{c}|=r,\qquad\forall c,$ (11)
where all class means have the same Euclidean norm $r$.
@@
Class-Mean Optimization. Under the neural collapse theory, feature embeddings are prepresented at the level of class means rather than individual samples. Training with a fixed classifier $W$ amounts to minimizing the cross-entropy over class means $\mu_{c}$:
$\min_{\{\mu_{c}\}}\ \ \mathbb{E}_{c\sim\pi}\!\left[-\log\frac{\exp\big(\gamma\,w_{c}^{\top}\mu_{c}\big)}{\sum_{k=1}^{C}\exp\big(\gamma\,w_{k}^{\top}\mu_{c}\big)}\right]\quad\text{s.t.}\quad\ \$(12)
where, $\mathbb{E}_{c\sim\pi}$ is the class index and $c$ is treated as a random variable according to the class probability $\pi$. $w_{k}$ is the weight vector of class $k$ and $w_{c}$ is the $c$-th column of $W$ for the target class. Also, with the optional constraint (11).
Margin Classification. The objective in 12 admits an equivalent max-margin formulation. Therefore, the classification margin [23] for class $c$ against class $k$ is expressed in a margin form:
$\max_{\{\mu_{c}\}}\ \ \min_{c\neq k}\ \big(w_{c}^{\top}\mu_{c}-w_{k}^{\top}\mu_{c}\big)\quad\text{s.t.}\quad,\ \$(13)
$\|\mu_{c}\|=r.$ (14)
@@
Relation to simplex-ETF. In standard learnable classifiers, neural collapse theory predicts that classifier weights align with class means, i.e., $w_{c}\propto\mu_{c}$. The centered and normalized class means form a simplex equiangular tight frame (ETF) $\langle\mu_{c},\mu_{c^{\prime}}\rangle$ [33], [16]:
@@
$\left\langle\mu_{c},\mu_{c^{\prime}}\right\rangle=\begin{cases}1,&c=c^{\prime},\\[4.0pt] -\dfrac{1}{C-1},&c\neq c^{\prime}.\end{cases}$ (15)
@@
However, in the FNOClassifier, $W$ is fixed with orthonormal, non-negative columns [8], satisfying (5). As a result, optimization of (12) does not learn the classifier directions, but instead aligns with class means $\mu_{c}$ with the fixed directions $w_{c}$ which are subject to zero-mean and equal-norm restrictions yielding a collapsed, balanced configuration without learning $W$.

### v3

@@ 3.1.2 Parallel Vision Mamba Module
The PVM module [26, 29], as shown on the left side of Fig. 4 (part a), mainly consists of Mamba integrated with residual connections, which improve Mamba’s ability to capture deep spatial relationships. Initially, the input $Y_{in}^{C}$ with channel $C$ goes through layer normalization. The features that are fed into Mamba are first branched out as $X_{1}^{C/4}$, $X_{2}^{C/4}$, $X_{3}^{C/4}$ and $X_{4}^{C/4}$, each with channel $C/4$. The Mamba outputs combined with the residual connection from the inputs and adjustment factor for optimization are then concatenated together to obtain the four feature maps. The right side of Fig. 4 (part b) shows the Mamba component used in the PVM layer. These features are concatenated to obtain $Y_{out}$ with the number of channels as $C$. The concatenated feature outputs are layer normalized and then projected using a Projection operation. By processing deep features in a parallel manner, the PVM module is able to capture multi-scale and intricate features using different kernel sizes. This process significantly reduces the parameters by retaining the same receptive field, which addresses the explosive influence of channel numbers since the number of parameters in Mamba is highly dependent on the number of input channels [29]. The specific operations are expressed by the following equations:
$X_{i}^{C/4}=S_{p}[\text{LN}(Y_{in}^{C})],\quad i=1,2,3,4$ (1)
$VM\_X_{i}^{C/4}$ $=Mamba(X_{i}^{C/4})+\theta\cdot X_{i}^{C/4},\quad i=1,2,3,4$ (2)
$Y_{\text{out}}=Cat\big(VM\_X_{1}^{C/4},\,VM\_X_{2}^{C/4},\,VM\_X_{3}^{C/4},\,VM\_X_{1}^{C/4}\big)$ (3)
$\text{Out}=Proj\big[\text{LN}(Y_{\text{out}})\big]$ (4)
Here, $LN$ denotes the layer normalization, $S_{p}$ is the split operation that splits the input into four separate branches, $\theta$ is the residual scaling factor, catcat denotes the concatenation operation where the outputs are concatenated and $Proj$ is the projection operation where it projects concatenated representation back into the feature space. From Eq. 2, we employed parallel Vision Mamba feature processing while keeping the total number of channels constant, allowing us to preserve high accuracy while achieving significant parameter reduction.
@@ 3.1.4 Fixed Non-Negative Orthogonal Classifier for Model Parameter Reduction
The Fixed Non-Negative Orthogonal Classifier or FNOClassifier [8] module is a lightweight neural classifier designed to map feature embeddings into class logits through structured and normalized projections. During initialization, the model partitions the input feature dimensions into approximately equal groups corresponding to the number of output classes. Random permutation ensures that each class is assigned a unique subset of features, with any remaining dimensions distributed randomly among the classes to maintain balance. A binary base matrix is then constructed, where each row indicates the feature subset associated with a class, and is subsequently L2-normalized along the row dimension to enforce orthogonality and numerical stability. The model also includes an optional input feature normalization step where input vectors are scaled by their L2 norm, ensuring consistent magnitude across samples. In the forward pass, the input which is either normalized or raw is linearly projected using the scaled base matrix, modulated by a learnable scalar parameter $W$ to produce class scores. This allows the FNOClassifier to serve as an efficient and interpretable projection-based classifier, emphasizing structured feature grouping and normalization for robust classification.
Fixed classifiers that incorporate orthogonality are recognized for their cost efficiency and their ability to sometimes outperform learnable classifiers on popular benchmarks. However, existing fixed orthogonal classifiers suffer from geometric limitations that fundamentally prevent them from invoking neural collapse [8, 30].
@@
The FNOClassifier was proposed to resolve the issue that previous fixed orthogonal classifiers fail to invoke neural collapse (NC). Neural collapse is a critical phenomenon in which last-layer features converge to the simplex ETF (Equiangular Tight Frame) structure necessary to achieve global optimality in a layer-peeled model, due to their inherent geometric limitations [8] [30]. The development of the FNOClassifier relies on an analysis of zero-mean NC that specifically considers orthogonality within non-negative Euclidean space. By satisfying the necessary properties of zero-mean NC, the FNOClassifier is designed to induce the optimal solution and maximize the margin of an orthogonal layer-peeled model. This classifier yields an inherent feature dimension separation effect by mitigating feature interference in continual learning and tackling the limitations of mixup on the hypersphere in imbalanced learning, ultimately demonstrating significant performance improvements across various experiments.
Fixed Non-Negative Orthogonal Classifier (FNOClassifier). Let $D$ and $F$ be the number of classes and feature dimension, respectively. A classifier with partial orthogonal weight matrix $W$ is employed as $W\in\mathbb{R}^{F\times D}$, where the weights are not learned during training and require
$W^{\top}W$ $=I_{D},\qquad W_{j,k}\geq 0\ \ (\text{element\mbox{-}wise}),$ (5)
where $I_{D}$ is identity matrix and $W_{j,k}$ is the $(j,k)$-th element of $W$. The orthogonality constraint ensures a distinct and independent direction for each class in the feature space, which maximizes the separation between classes and minimizes classifier-induced bias. The non-negativity constraint limits the directions of the classifier to the positive orthant promoting stable and consistent alignment between trained feature representations and their corresponding class-vectors.
Based on an input feature representation, the logits produced by the classifier are defined by
@@
$\qquad p(d\mid x)=\mathrm{softmax}_{d}\big(z(x)\big).$ (7)
where $z(x)$ is the logit vector, $\gamma$ is a positive scaling parameter controlling the magnitude of the logits and $x$ is the input feature representation. $p(d|x)$ is the predicted probability of an input sample with feature representation where $x$ belongs to class $d$.
Neural collapse. Neural collapse refers to a set of geometric properties that emerge toward the end phase of training neural networks using crossentropy loss. First, intra-class collapse occurs. Under zero-mean neural collapse, features from the same class converge toward a shared class mean where $x_{l}^{(d)}$ is the feature embedding of the $l$-th sample belonging to the class $d$, $n_{d}$ and $\mu_{d}$ are the number of samples and the class mean feature vector for class $d$, respectively and defined as:
$\mu_{d}=\frac{1}{n_{d}}\sum_{l=1}^{n_{d}}x_{l}^{(d)}$ (8)
The embeddings satisfy (9) as the training progresses
$x_{l}^{(d)}$ $\xrightarrow[]{\text{training}}\mu_{d}\quad\text{(intra-class collapse)}$ (9)
@@
Second, class mean centering occurs. Under zero-mean neural collapse, the weighted mean of the class becomes centered and converges to zero or origin [16], [7], [2]. Therefore, the features behave as:
$\sum_{d=1}^{D}\psi_{d}\,\mu_{d}$ $=0\quad\text{(centered class means)}.$ (10)
@@
In (10), $D$ is the total number of classes, the class probability is denoted by $\psi_{d}$ where the global feature distribution is centered at origin and the weighted average of all class means is zero, thus, ensuring a balanced and symmetric arrangement of class means in the feature space.
Finally, neural collapse enforces the equal-norm constraint on class means, which is given as:
$\|\mu_{d}\|=r,\qquad\forall_{d},$ (11)
where all class means have the same Euclidean norm $r$. This constraint ensures equal margins across classes in the feature space by preventing dominance through the larger feature magnitudes. Integrated with intra-class collapse and centered class means, the resulting highly structured feature geometry concentrate samples from each class tightly around their respective means where the class means are evenly distributed around the origin and the angles between class directions are maximized. This yields compact, balanced, and maximally separated class representations consistent with Equiangular Tight Frame (ETF) geometry [16], [33].
@@
Class-Mean Optimization. Under the neural collapse theory, feature embeddings are represented at the level of class means rather than individual samples. Optimization of class mean representations $\mu_{d}$ becomes equivalent to optimization of individual feature vectors as intra-class variance diminishes. Consequently, training with a fixed classifier $W$ amounts to minimizing the cross-entropy over class means $\mu_{d}$:
$\min_{\{\mu_{d}\}}\ \ \mathbb{E}_{d\sim\psi}\!\left[-\log\frac{\exp\big(\gamma\,w_{d}^{\top}\mu_{d}\big)}{\sum_{m=1}^{D}\exp\big(\gamma\,w_{m}^{\top}\mu_{d}\big)}\right]\quad\text{s.t.}\quad\|\mu_{d}\|=r\quad\ \$(12)
where, $d$ is the class index treated as a random variable obtained from the class prior $\psi$ and $\mathbb{E}_{d\sim\psi}[\cdot]$ denotes the average over $d$ under the class distribution $\psi$. $w_{m}$ is the weight vector of the competing class $m$ and $w_{d}$ is the $d$-th column of $W$ for the target class, where $m\neq d$. Under this formulation, optimization solely acts on the feature extractor, aligning each class mean $\mu_{d}$ in the direction of its corresponding classifier while maintaining the separation from the directions associated with other classes. When integrated with equal-norm constraint in (11), it promotes a balanced and symmetric feature geometry consistent with ETF structure of neural collapse [16].
Margin Classification. The objective in 12 is closely related to an implicit max-margin perspective. In certain overparameterized models, training with cross-entropy loss implicitly maximizes the classification margin between classes [23]. Therefore, the classification margin [23] for class $d$ versus class $m$ is expressed in the following margin form:
$\max_{\{\mu_{d}\}}\ \ [\min_{d\neq m}\ \big(w_{d}^{\top}\mu_{d}-w_{m}^{\top}\mu_{d}\big)]\quad\text{s.t.}\quad\|\mu_{d}\|=r\quad$ (13)
where the inner minimization ensures separation from the closest competing class. This drives each class mean $\mu_{d}$ to align with its associated classifier direction $w_{d}$ while maintaining the distance from other class directions. The equal-norm constraint ensures min-margin, max-margin decision boundaries consistent with neural collapse.
@@
Relation to simplex-ETF. In standard learnable classifiers, neural collapse theory predicts that classifier weights align with class means, i.e., $w_{d}\propto\mu_{d}$. The centered and normalized class means form a simplex equiangular tight frame (ETF) $\langle\tilde{\mu}_{d},\tilde{\mu}_{m}\rangle$ $\forall$ $m,d\in\{1,...,D\}$ [33], [16]:
@@
$\left\langle\tilde{\mu}_{d},\tilde{\mu}_{m}\right\rangle=\begin{cases}1,&d=m,\\[4.0pt] -\dfrac{1}{D-1},&d\neq m.\end{cases}$ (14)
@@
However, in FNOClassifier, $W$ is fixed with orthonormal, non-negative columns [8], satisfying (5). As a result, optimization of (12) does not learn the classifier directions, but instead aligns the class means $\mu_{d}$ with the fixed directions $w_{d}$ which are subject to zero-mean and equal-norm restrictions, producing a collapsed, balanced configuration without learning the partial orthogonal weight matrix $W$.